OpenStudy (sh3lsh):

Decompose into partial fractions: 12/x^2-9

OpenStudy (anonymous):

2/ (x -3) - 2/ (x + 3)

OpenStudy (lgbasallote):

that\s 12/(x+3)(x-3) right? so your partial fraction is A/x+3 + B/x-3 correct? so 12/(x+3)(X-3) = a/(X+3) + B/x-3 multiply both sides by (x+3)(x-3)...the LCD that gives you 12 = A(x-3) + B(x+3) expand... 12 = Ax - 3A + Bx +3B now use gaussian method (if you're familiar with it) COEFFICIENTS OF x 0 = A + B COEFFICIENT OF CONSTANT 12 = -3a + 3B 4 = B -A system of linear eq... B-A = 4 B + A = 0 2b =4 b = 2 so a = -2 so your equation is 2/x+3 - 2/x-3 do you understand?

OpenStudy (sh3lsh):

wow. Yes. Thanks so much lgbasallote!!!